# Coupling without Communication and Drafter-Invariant Speculative Decoding

### v2

@@ 1.3 Related Work
Most closely related to our work are prior results on weighted coordinated sampling methods, which generalize well-known techniques like MinHash and the $k$-minimum values sketch for unweighted coordinated sampling (Broder, 1997; Broder et al., 1998; Cohen and Kaplan, 2007; Beyer et al., 2007; Li et al., 2006). The goal in weighted coordinated sampling is quite similar to ours: Alice and Bob hold vectors, $A$ and $B$, and seek to independently produce a subsample of entries from their vector so that 1) indices corresponding to larger entries in the vectors are sampled with higher probability and 2) Alice and Bob return many of the same indices. The aim is to use Alice and Bob’s subsamples to estimate functions involving interactions between corresponding entries in $A$ and $B$, like the inner product $\langle A,B\rangle=\sum_{i=1}^{n}A_{i}B_{i}$ (Daliri et al., 2024a; Li, 2017). To do so effectively, it is critical that we have access to pairs $(A_{i},B_{i})$ with the same index $i$.
Weighted coordinated sampling methods include Weighted MinHash (also referred to as “consistent weighted sampling”) (Manasse et al., 2010; Ioffe, 2010; Haeupler et al., 2014; Wu et al., 2020), Gumbel sampling (Cohen, 2015, 2023), threshold sampling (Flajolet, 1990; Duffield et al., 2005), priority sampling (Duffield et al., 2004; Daliri et al., 2024b), and more (Estan and Naughton, 2006). While all of these methods intuitively seek to generate samples from $A$ and $B$ that contain the same indices with high probability, that is typically not the final goal: the methods are analyzed for specific downstream application. As such, we are unaware of prior work that specifically obtains bounds for our “communication-free coupling” problem.
We note that several coordinated sampling methods cannot be applied in our context. For example, priority sampling does not ensure that entries from $A$ and $B$ are sampled with probability exactly proportional to specified probabilities (we require samples to truly be drawn from $\mathcal{P}$ and $\mathcal{Q}$), and threshold sampling does not return a fixed number of samples (we always need exactly one sample). Of the major methods, this leaves Weighted MinHash and Gumbel sampling, which are the two methods we analyze.
@@ 2 Preliminaries
Notation. Throughout the paper, we assumes two parties, Alice and Bob, who hold discrete distributions $\mathcal{P}$ and $\mathcal{Q}$ over the set $\{1,\ldots,n\}$. We use $p_{i}$ to denote the probability that $x$ drawn from $\mathcal{P}$ equals $i$ and $q_{i}$ to denote the probability that $x$ drawn from $\mathcal{Q}$ equals $i$. We will sometimes write $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ and $\mathcal{Q}=\{q_{1},\ldots,q_{n}\}$ since the list of probabilities completely specifies the distributions.
@@
1:Await for $(a,\mathcal{P})$ from Alice.
2:With probability $\min(1,q_{a}/p_{a})$ set $b=a$. Otherwise, sample $b\sim\mathcal{Q}^{\prime}$, where $\mathcal{Q}^{\prime}=\{q_{1}^{\prime},\ldots,q_{n}^{\prime}\}$ is a distribution on $\{1,\ldots,n\}$ with $q_{i}^{\prime}=\frac{\max(0,q_{i}-p_{i})}{\sum_{j=1}^{n}\max(0,q_{j}-p_{j})}$.
Clearly Protocol 1 ensures that $a\sim\mathcal{P}$. It is easily checked that it also ensures that $b\sim\mathcal{Q}$. In particular, if $q_{i}\leq p_{i}$, $\max(0,q_{i}-p_{i})=0$, so we only set $b=i$ in step one of the protocol. We do so with probability $\Pr[a=i]\cdot q_{i}/p_{i}=q_{i}$, as desired. If $q_{i}\geq p_{i}$, then with $p_{i}$ probability, we set $b=i$ in step one of the protocol (exactly when $a$ is set to $i$). There is also some chance we set $b=i$ in the second step of the protocol, which we execute with probability $\sum_{j=1}^{n}p_{j}\cdot(1-\min(1,q_{j}/p_{j}))=\sum_{j=1}^{n}\max(0,p_{j}-q_{j})=\sum_{j=1}^{n}\max(0,q_{j}-p_{j})$. So, the overall probability we set $b=i$ in the second step is $q_{i}^{\prime}\cdot\sum_{j=1}^{n}\max(0,q_{j}-p_{j})=q_{i}-p_{i}$. Thus, the total probability we set $b=i$ (in either step) is $q_{i}-p_{i}+p_{i}=q_{i}$, as desired.
@@
3 Communication-free Protocols
We begin by describing two communication-free protocols that achieve the bound of 1, so nearly match the optimal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ bound obtainable (with communication) by Protocol 1.
@@ 3.1 Weighted MinHash
We first consider a method based on the Weighted MinHash algorithm, which was introduced under the name “consistent weighted sampling” by Manasse et al. (2010). The fact that this method achieves the bound of 1 was observed in prior work, including by Manasse et al. (2010). We provide a self-contained analysis for completeness, and to obtain a sharp bound (i.e., an exact expression for $\Pr[a=b]$).
There are many different ways of implementing the Weighted MinHash method. See Wu et al. (2020) for an overview. We analyze a particularly simple implementation suggested by Shrivastava (2016), which is detailed in Protocol algorithm 2 and illustrated in Figure 1.
@@
Figure 1: The Weighted MinHash method, described in Protocol 2, selects a sample from a distributions on $n$ items with probabilities $p_{1},\ldots,p_{n}$ by drawing a sequence of numbers $u_{1},u_{2},\ldots,$ uniformly at random from the interval $[0,n]$. The first time one of these numbers lands in a subinterval $[j-1,j-1+p_{j}]$ for any $j\in 1,\ldots,n$ , the protocol returns $j$. For example, in the illustration above, Alice returns $2$ (since $u_{2}$ falls within $[1,1+p_{2}]$) and Bob returns $4$ (since $u_{5}$ falls within $[3,3+p_{4}]$)
@@
As illustrated in Figure 1, to run Protocol 2, Alice draws values uniformly at random from the interval $[0,n]$. The protocol terminates and returns $j$ as soon as a value lands in any subinterval $[j-1,j-1+p_{j}]$, which are shaded in the top row of Figure 1. Since these subintervals have lengths $p_{1},\ldots,p_{n}$, the probability that Alice returns $j$ is exactly $\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}=p_{j}$. The proof that Bob returns $j$ with probability $q_{j}$ is identical.
@@
If $u_{k}$ does land in the interval $[j-1+\min(p_{j},q_{j}),j-1+\max(p_{j},q_{j})]$, Alice is the party who terminates and return $j$ whenever $p_{j}>q_{j}$. Otherwise, if $q_{j}>p_{j}$, Bob terminates and returns $j$. So, the question becomes, with what probability with the other party also returns $j$ on a future iteration? Future iterations involve random variables completely independent from $u_{k}$, so the probability is simply $\min(p_{j},q_{j})$.
@@
3.2 Gumbel Sampling
The next method we analyze is based on Gumbel sampling, an existing protocol that is widely for sampling from discrete distributions, even when coupling is not necessary. Under various names like PPSWOR, order sampling, and bottom-k sampling, the method has found widespread application across randomized algorithms (Rosén, 1997; Cohen, 2023; Efraimidis and Spirakis, 2006). We give pseudocode in Protocol 3.
@@ 3.2 Gumbel Sampling
Our main result on Gumbel sampling is Theorem 2, which we restate below: See 2
To prove Theorem 2, we require some standard preliminaries about exponential random variables. For a reference, see, e.g. Cohen (2023).
@@
A random variables $X\sim\exp(\lambda)$ is exponential with parameter $\lambda$ if $x=-\ln(u)/\lambda$ for a random variable $u$ drawn uniformly from $[0,1]$. Suppose we have independent exponential random variables $X_{1}\sim\exp(\lambda_{1})$ and $X_{2}\sim\exp(\lambda_{2})$. Then we have:
@@
So, we focus on analyzing $\Pr[a=j\text{ and }b=j]$ for a fixed $j$. We have that $a=j$ and $b=j$ if for all $i\neq j$,
@@
Observe that $-\ln(u_{j})$ is an exponential random variable with parameter $\lambda=1$ and, applying 6, $\min_{i\neq j}\left[-\ln(u_{i})\cdot\frac{1}{\max(p_{i}/p_{j},q_{i}/q_{j})}\right]$ is exponential with parameter $\lambda=\sum_{i\neq j}\max(p_{i}/p_{j},q_{i}/q_{j})$. These two random variables are independent, so again by 6:
@@
To simplify this expression, assume for now that $p_{j}\leq q_{j}$. We can make a symmetric argument for the case when $q_{j}>p_{j}$. If $p_{j}\leq q_{j}$, we have:
@@
4 Optimality of Communication-free Protocols
In the previous section, we showed that two protocols, one based on Weighted MinHash, and the other on Gumbel Sampling, allow Alice and Bob to sample $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ without communication in such a way that $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$. A natural question is if better protocols exist, especially given that we know we can achieve a bound of $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ when communication is allowed. In this section, we show that this is impossible in the worst-case. In particular, we prove the following result, restated from Section 1:
See 3
Proof.
Now, suppose a protocol is used to sample $a\sim\mathcal{P},b\sim\mathcal{Q}$ and $c\sim\mathcal{R}$. For any $i\in\{1,2,3\}$, it must be that one of $a,b$ or $c$ does not equal $i$, since at least one of the distributions has 00 mass on item $i$. It follows that only one of the equalities $a=b$, $a=c$, and $b=c$ holds at once. So, we have that for any protocol, $\min(\Pr[a=b],\Pr[a=c],\Pr[b=c])\leq 1/3$. I.e., for two of the distributions, the protocol gives collision probability $\leq 1/3$. The result then follows from simply verifying that that $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})=D_{\mathrm{TV}}(\mathcal{P},\mathcal{R})=D_{\mathrm{TV}}(\mathcal{Q},\mathcal{R})=1/2$:
We can generalize this lower bound by having $d+1$ distributions $\mathcal{P}_{1},\ldots,\mathcal{P}_{d+1}$ over $11,\ldots,d+$ 1. Specifically, let the $i^{\text{th}}$ distribution, $\mathcal{P}_{i}$, be uniform over all $j\neq i$. So, $x\sim\mathcal{P}_{i}$ has $1/d$ probability of equaling any $j\neq i$. It can be checked that $D_{\mathrm{TV}}(\mathcal{P}_{i},\mathcal{P}_{j})=1/d$ for all $i,j$.
Now, consider a protocol that samples $x_{1}\sim\mathcal{P}_{1},\ldots,x_{d+1}\sim\mathcal{P}_{d+1}$. Consider all (${d+1\choose 2}$ equalities of the form $\{x_{i}=x_{j},i\neq j\}$. No matter what the outcome of $x_{1},\ldots,x_{d+1}$, we claim that at least $d$ of these inequalities will not be satisfied. In particular, as in the argument above, it cannot be that $x_{1},\ldots,x_{d}$ all have the same value $v$ since one of $\mathcal{P}_{1},\ldots,\mathcal{P}_{d+1}$ has zero mass for any possible $v$. As such, at least two items in $x_{1},\ldots,x_{d+1}$ must be different from each other. Suppose without loss of generality that $x_{1}\neq x_{2}$.
First, the equality $x_{1}=x_{2}$ is immediately not satisfied. Then for all $j\neq 1,2$, it must be that either $x_{j}\neq x_{1}$ or $x_{j}\neq x_{2}$, which is $d+1-2=d-1$ additional inequalities not satisfied. So, no matter what the outcome of the protocol, the ratio of equalities of the form $\{x_{i}=x_{j},i\neq j\}$ that are satisfied is at most:
$1-\frac{d}{{d+1\choose 2}}=1-\frac{2}{d+1}.$
It follows that there must be some pair $i,j$ such that:
$\Pr[x_{i}=x_{j}]\leq 1-\frac{2}{d+1}=\frac{1-1/d}{1+1/d}.$
We remark that, in spite of Theorem 3, both Weighted MinHash and Gumbel sampling often perform better in practice than the worst-case bound. For example, in our experiments on speculative decoding in Section 5, Gumbel sampling often nearly matches optimal coupling, obtaining collision probability close to $1-\mathcal{D}_{TV}(\mathcal{P},\mathcal{Q})$. Exploring this observation further is an interesting direction for future work.
It maybe be that the bound of Theorem 2 tends to be much stronger for the sort of “quickly decaying” distributions common in applications. As an extreme example, suppose that $n=2$, or equivalently, $\mathcal{P}$ and $\mathcal{Q}$ only place mass on two items. Then Gumbel sampling has the same collision probability as that of optimal coupling. To see this, without loss of generality, let $p_{1}\leq q_{1}$, then the result of Theorem 2 simplifies as:
$\Pr[a=b]=\frac{1}{1+\max\left(\frac{p_{1}}{p_{2}},\frac{q_{1}}{q_{2}}\right)}+\frac{1}{1+\max\left(\frac{p_{2}}{p_{1}},\frac{q_{1}}{q_{1}}\right)}=\frac{1}{1+\frac{q_{1}}{q_{2}}}+\frac{1}{1+\frac{p_{2}}{p_{1}}}=q_{2}+p_{1}=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}).$

### v4

@@ 1.3 Related Work
Our work is closely related to results on weighted coordinated sampling methods, which generalize well-known techniques like MinHash and the $k$-minimum values sketch for unweighted coordinated sampling (Broder, 1997; Broder et al., 1998; Cohen and Kaplan, 2007; Beyer et al., 2007; Li et al., 2006). The goal in weighted coordinated sampling is similar to ours: Alice and Bob hold vectors, $A$ and $B$, and seek to independently produce a subsample of entries from their vector so that 1) indices corresponding to larger entries in the vectors are sampled with higher probability and 2) Alice and Bob return many of the same indices. The aim is to use Alice and Bob’s subsamples to estimate functions involving interactions between corresponding entries in $A$ and $B$, like the inner product $\langle A,B\rangle=\sum_{i=1}^{n}A_{i}B_{i}$ (Li, 2017; Bessa et al., 2023; Daliri et al., 2024a). To do so effectively, it is critical that we have access to pairs $(A_{i},B_{i})$ with the same index $i$.
Weighted coordinated sampling methods include Weighted MinHash (also referred to as “consistent weighted sampling”) (Manasse et al., 2010; Ioffe, 2010; Haeupler et al., 2014; Wu et al., 2020), Gumbel sampling (Cohen, 2015, 2023), threshold sampling (Flajolet, 1990; Duffield et al., 2005), priority sampling (Duffield et al., 2004; Daliri et al., 2024b), and more (Estan and Naughton, 2006). While all of these methods intuitively seek to generate samples from $A$ and $B$ that contain the same indices with high probability, that is typically not the final goal: the methods are analyzed for specific downstream applications. As such we are unaware of any prior work besides that of Bavarian et al. (2020) which specifically obtains bounds for our “communication-free coupling” problem.
We also note that several coordinated sampling methods cannot be applied in our context. For example, priority sampling does not ensure that entries from $A$ and $B$ are sampled with probability exactly proportional to specified probabilities (we require samples to truly be drawn from $\mathcal{P}$ and $\mathcal{Q}$), and threshold sampling does not return a fixed number of samples (we always need exactly one sample). Of the major methods, this leaves Weighted MinHash and Gumbel sampling, which are the two methods we address in this paper.
@@ 2 Preliminaries
Notation. Throughout the paper, we assume two parties, Alice and Bob, who hold discrete distributions $\mathcal{P}$ and $\mathcal{Q}$ over the set $\{1,\ldots,n\}$. We use $p_{i}$ to denote the probability that $x$ drawn from $\mathcal{P}$ equals $i$ and $q_{i}$ to denote the probability that $x$ drawn from $\mathcal{Q}$ equals $i$. We will sometimes write $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ and $\mathcal{Q}=\{q_{1},\ldots,q_{n}\}$ since the list of probabilities completely specifies the distributions.
@@
1:Await for $(a,\mathcal{P})$ from Alice. With probability $\min(1,q_{a}/p_{a})$ set $b=a$.
2:Otherwise, sample $b\sim\mathcal{Q}^{\prime}$, where $\mathcal{Q}^{\prime}=\{q_{1}^{\prime},\ldots,q_{n}^{\prime}\}$ is a distribution with $q_{i}^{\prime}=\frac{\max(0,q_{i}-p_{i})}{\sum_{j=1}^{n}\max(0,q_{j}-p_{j})}$.
Clearly Protocol 1 ensures that $a\sim\mathcal{P}$. It is easily checked that it also ensures that $b\sim\mathcal{Q}$. In particular, for any $i$ where $q_{i}\leq p_{i}$, $\max(0,q_{i}-p_{i})=0$, so we can only set $b=i$ in Step 1 of the protocol. We do so with probability $\Pr[a=i]\cdot q_{i}/p_{i}=q_{i}$, as desired. If $q_{i}\geq p_{i}$, then with probability $p_{i}$, we set $b=i$ in Step 1 of the protocol (exactly when $a$ is set to $i$). There is also some chance we set $b=i$ in Step 2 of the protocol, which we execute with probability $\sum_{j=1}^{n}p_{j}\cdot(1-\min(1,q_{j}/p_{j}))=\sum_{j=1}^{n}\max(0,p_{j}-q_{j})=\sum_{j=1}^{n}\max(0,q_{j}-p_{j})$. So, the overall probability we set $b=i$ in the second step is $q_{i}^{\prime}\cdot\sum_{j=1}^{n}\max(0,q_{j}-p_{j})=q_{i}-p_{i}$. Thus, the total probability we set $b=i$ (in either step) is $q_{i}-p_{i}+p_{i}=q_{i}$, as desired.
@@
3 An Existing Communication-free Protocol
As discussed, a communication-free protocol based on the Weighted MinHash algorithm can nearly match the performance of Protocol 1 (which requires communication, since Bob needs to know Alice’s distribution $\mathcal{P}$). A number of papers analyze Weighted MinHash, including Bavarian et al. (2020). We provide a self-contained analysis below for completeness, and to obtain a sharp bound (i.e., an exact expression for $\Pr[a=b]$).
@@ 3.1 Weighted MinHash
Technically speaking, there are many different ways to implement Weighted MinHash, which might be more accurately described as a family of closely-related sampling method. See (Wu et al., 2020) for an overview. We analyze a particularly simple implementation suggested by a number of authors (Kleinberg and Tardos, 2002; Holenstein, 2009; Shrivastava, 2016). It is detailed in Protocol 2 and illustrated in Figure 1.
@@
Figure 1: The Weighted MinHash method, described in Protocol 2, selects a sample from a distributions on $n$ items with probabilities $p_{1},\ldots,p_{n}$ by drawing a sequence of numbers $u_{1},u_{2},\ldots,$ uniformly at random from the interval $[0,n]$. The first time one of these numbers lands in a subinterval $j-1,j-1+p_{j}]$ for any $j\in 1,\ldots,n$ , the protocol returns $j$. For example, in the illustration above, Alice returns 22 (since $u_{2}$ falls within $1,1+p_{2}]$) and Bob returns 44 (since $u_{5}$ falls within $3,3+p_{4}]$)
@@
As illustrated in Figure 1, to run Protocol 2, Alice draws values uniformly at random from the interval $[0,n]$. The protocol terminates and returns $j$ as soon as a value lands in any subinterval $j-1,j-1+p_{j}]$, which are shaded in the top row of Figure 1. Since these subintervals have lengths $p_{1},\ldots,p_{n}$, the probability that Alice returns $j$ is exactly $\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}=p_{j}$. The proof that Bob returns $j$ with probability $q_{j}$ is identical.
@@
If $u_{k}$ does land in the interval $j-1+\min(p_{j},q_{j}),j-1+\max(p_{j},q_{j})]$, Alice is the party who terminates and return $j$ whenever $p_{j}>q_{j}$. Otherwise, if $q_{j}>p_{j}$, Bob terminates and returns $j$. So, the question becomes, with what probability does the other party also returns $j$ on a future iteration? Future iterations involve random variables completely independent from $u_{k}$, so the probability is simply $\min(p_{j},q_{j})$.
@@
3.2 Optimality of Communication-free Protocols
As discussed in Section 1, when parameterizing performance by total variance distance, it can be proven that the bound of 5 for Weighted MinHash is optimal for worst-case distributions. Bavarian et al. (2020) do so via a reduction from an easier “constrained agreement problem”. We provide an even simpler proof here, first restating the formal optimality claim from Section 1:
See 2
Proof.
Now, suppose a protocol is used to sample $a\sim\mathcal{P},b\sim\mathcal{Q}$ and $c\sim\mathcal{R}$. For any $i\in\{1,2,3\}$, it must be that one of $a,b$ or $c$ does not equal $i$, since at least one of the distributions has 0 mass on item $i$. It follows that only one of the equalities $a=b$, $a=c$, and $b=c$ holds at once. So, we have that for any protocol, $\min(\Pr[a=b],\Pr[a=c],\Pr[b=c])\leq 1/3$. I.e., for two of the distributions, the protocol gives collision probability $\leq 1/3$. The result then follows from simply verifying that that $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})=D_{\mathrm{TV}}(\mathcal{P},\mathcal{R})=D_{\mathrm{TV}}(\mathcal{Q},\mathcal{R})=1/2$:
We can generalize this result by having $d+1$ distributions $\mathcal{P}_{1},\ldots,\mathcal{P}_{d+1}$ over $1,\ldots,d+1$. Specifically, let the $i^{\text{th}}$ distribution, $\mathcal{P}_{i}$, be uniform over all $j\neq i$. So, $x\sim\mathcal{P}_{i}$ has $1/d$ probability of equaling any $j\neq i$. It can be checked that $D_{\mathrm{TV}}(\mathcal{P}_{i},\mathcal{P}_{j})=1/d$ for all $i,j$.
Now, consider a protocol that samples $x_{1}\sim\mathcal{P}_{1},\ldots,x_{d+1}\sim\mathcal{P}_{d+1}$. Consider all $d+1\choose 2$} equalities of the form $\{x_{i}=x_{j},i\neq j\}$. No matter what the outcome of $x_{1},\ldots,x_{d+1}$, we claim that at least $d$ of these inequalities will not be satisfied. In particular, as in the argument above, it cannot be that $x_{1},\ldots,x_{d}$ all have the same value $v$ since one of $\mathcal{P}_{1},\ldots,\mathcal{P}_{d+1}$ has zero mass for any possible $v$. As such, at least two items in $x_{1},\ldots,x_{d+1}$ must be different from each other. Suppose without loss of generality that $x_{1}\neq x_{2}$.
First, the equality $x_{1}=x_{2}$ is immediately not satisfied. Then for all $j\neq 1,2$, it must be that either $x_{j}\neq x_{1}$ or $x_{j}\neq x_{2}$, which is d+12=d1d+1-2=d-1 additional inequalities not satisfied. So, no matter what the outcome of the protocol, the ratio of equalities of the form $\{x_{i}=x_{j},i\neq j\}$ that are satisfied is at most:
$1-\frac{d}{{d+1\choose 2}}=1-\frac{2}{d+1}.$
It follows that there must be some pair $i,j$ such that:
$\Pr[x_{i}=x_{j}]\leq 1-\frac{2}{d+1}=\frac{1-1/d}{1+1/d}.$
4 Better Coupling via Gumbel Sampling
Despite the limitation of LABEL:{clm:ub}, in this section we show that it is nevertheless possible to strictly improve on the Weighted MinHash method from the previous section. To do so, we analyze a method based on Gumbel sampling, an existing protocol that is widely used for sampling from discrete distributions, even when coupling is not necessary. Under various names like PPSWOR, order sampling, and bottom-k sampling, the method has found widespread application across randomized algorithms (Rosén, 1997; Cohen, 2023; Efraimidis and Spirakis, 2006). We give pseudocode in Protocol 3.
@@ 4 Better Coupling via Gumbel Sampling
Our main result on Gumbel sampling is Theorem 3, which we restate below: See 3
As discussed in Section 1, it is not hard to find distributions for which the inequalities above are strict. In fact, Gumbel sampling often obtains a collision probability to the optimal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Understanding this strong practical performance is an interesting direction for future work. It may be be that the bound of Theorem 3 tends to be much stronger for the sort of “quickly decaying” distributions common in applications. As an extreme example, suppose that $n=2$, or equivalently, $\mathcal{P}$ and $\mathcal{Q}$ only place mass on two items. Then, as above, Gumbel sampling has exactly the same collision probability as that of the optimal coupling. To see this, without loss of generality, let $p_{1}\leq q_{1}$, then the result of Theorem 3 simplifies as:
$\Pr[a=b]=\frac{1}{1+\max\left(\frac{p_{1}}{p_{2}},\frac{q_{1}}{q_{2}}\right)}+\frac{1}{1+\max\left(\frac{p_{2}}{p_{1}},\frac{q_{2}}{q_{1}}\right)}=\frac{1}{1+\frac{q_{1}}{q_{2}}}+\frac{1}{1+\frac{p_{2}}{p_{1}}}=q_{2}+p_{1}=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}).$
We prove Theorem 3 in two parts, first showing that $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$. We then prove that $\Pr[a=b]\geq C_{\text{WMH}}$ which is more difficult. To do so, we require some standard preliminaries about exponential random variables. For a reference, see, e.g. Cohen (2023).
@@
A random variable $X\sim\operatorname{Exp}(\lambda)$ is exponential with parameter $\lambda$ if $X=-\ln(u)/\lambda$ for a random variable $u$ drawn uniformly from $[0,1]$. Suppose we have independent exponential random variables $X_{1}\sim\operatorname{Exp}(\lambda_{1})$ and $X_{2}\sim\operatorname{Exp}(\lambda_{2})$. Then we have:
@@ 4.1 Main Analysis
So, we focus on analyzing $\Pr[a=j\text{ and }b=j]$ for a fixed $j$. Note that the probability is 0 is either $p_{j}=0$ or $q_{j}=0$, so we assume below that $\min(p_{j},q_{j})>0$. We have that $a=j$ and $b=j$ if for all $i\neq j$,
@@
Observe that $\ln(u_{j})$ is an exponential random variable with parameter $\lambda=1$ and, applying 6, $\min_{i\neq j}\left[-\ln(u_{i})\cdot\frac{1}{\max(p_{i}/p_{j},q_{i}/q_{j})}\right]$ is exponential with parameter $\lambda=\sum_{i\neq j}\max(p_{i}/p_{j},q_{i}/q_{j})$. These two random variables are independent, so again by 6:
@@
Summing over all $j$ for which $\min(p_{j},q_{j})>0$ yields the exactly expression for $\Pr[a=b]$ in theorem 3.
We next simplify this expression to prove that $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$, which establishes that Gumbel sampling at least matches the worst-case bound of Theorem 2. Assume for now that $p_{j}\leq q_{j}$. We can make a symmetric argument for the case when $q_{j}>p_{j}$. If $p_{j}\leq q_{j}$, we have:
@@
Pareto Improvement over Weighted MinHash
Next, we strengthen the above claim by showing that, in fact, $\Pr[a=b]\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$. I.e., that Gumbel sampling offers a Pareto improvement over Weighted MinHash.
To prove the stronger bound, we will analyze $\Pr[a=j\text{ and }b=j]$ separately for each $j\in 1,\ldots,n$. Again, we restrict to the case when $\min(p_{j},q_{j})>0$, since if either probability equals 0, then $\Pr[a=j\text{ and }b=j]$ equals 0 for both Gumbel sampling and Weighted MinHash. When $a$ and $b$ are sampled from $\mathcal{P}$ and $\mathcal{Q}$ via Weighted MinHash, it can be seen from the analysis in Section 3 that $\Pr[a=j\text{ and }b=j]={\min(p_{j},q_{j})\left(1+|p_{j}-q_{j}|\right)}/{\sum_{i=1}^{n}\max(p_{i},q_{i})}.$ So, based on our exact expression for the collision probability for Gumbel sampling from (7), it suffices to prove that:
$\frac{\min(p_{j},q_{j})\left(1+|p_{j}-q_{j}|\right)}{\sum_{i=1}^{n}\max(p_{i},q_{i})}\leq\frac{1}{\sum_{i=1}^{n}\max\left({p_{i}}/{p_{j}},{q_{i}}/{q_{j}}\right)}.$ (8)
Summing over all $j$ would then establish our desired bound.
To prove (8), we first note that, without loss of generality, we can assume $p_{j}\leq q_{j}$ and let $q_{j}=k\cdot p_{j}$ for some $k\geq 1$. Substituting and rearranging, we can then restate the (8) as follows:
$p_{j}\left(1+(k-1)p_{j}\right)\cdot\sum_{i=1}^{n}\max\left({p_{i}}/{p_{j}},{q_{i}}/{kp_{j}}\right)$ $\leq\sum_{i=1}^{n}\max(p_{i},q_{i}).$
To prove the above, it suffices to show that:
$\sum_{i=1}^{n}S_{i}$ $\geq 0,$ where $S_{i}$ $\mathbin{\stackrel{{\scriptstyle\rm def}}{{=}}}\left(\max(p_{i},q_{i})-\left(1+(k-1)p_{j}\right)\cdot\max\left(p_{i},{q_{i}}/{k}\right)\right).$ (9)
We analyze each term in the summation by splitting into two cases:
Case 1: $p_{i}>\frac{q_{i}}{k}$.
In this case, we have:
$S_{i}=\max(p_{i},q_{i})-\left(1+(k-1)p_{j}\right)p_{i}\geq p_{i}-\left(1+(k-1)p_{j}\right)p_{i}=(1-k)p_{j}p_{i}.$
Case 2: $p_{i}\leq\frac{q_{i}}{k}$.
In this case, we have:
$S_{i}=\left(1-\frac{1}{k}-(k-1)\frac{p_{j}}{k}\right)q_{i}=\frac{k-1}{k}(1-p_{j})q_{i}\geq(k-1)(1-p_{j})p_{i}.$
We conclude that:
$\sum_{i=1}^{n}S_{i}\geq(1-k)p_{j}\cdot\left[\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}\right]+(k-1)(1-p_{j})\cdot\left[\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}\right].$
Finally, recall that, by definition, $q_{j}=kp_{j}$. So the summation $\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}$ excludes the case $i=j$, and thus must be $\leq 1-p_{j}$. Similarly, $\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}\geq p_{j}$. Recalling that $k\geq 1$, so (1k)(1-k) is non-positive and (k1)(k-1) is non-negative, we conclude that:
$\sum_{i=1}^{n}S_{i}\geq(1-k)p_{j}(1-p_{j})+(k-1)(1-p_{j})p_{j}=0.$
We have thus proven equation (9) and (8). As discussed, summing (8) over all $j$ with $\min(p_{j},q_{j})>0$ establishes that, if $a,b$ are sampled from $\mathcal{P},\mathcal{Q}$ via Gumbel sampling, then $\Pr[a=b]\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$. ∎
Remark.
We conclude by observing that the main inequality $\Pr[a=b]\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$ in Theorem 3 is actually strict under very mild conditions, i.e., in most cases Gumbel sampling strictly improves on Weighted MinHash. In particular, consider the case when there are just three indices $j,k,l$ for which $p_{j}\neq q_{j}$, and $p_{k}\neq q_{k}$, and $p_{l}\neq q_{l}$, and one of the pairs $p_{l},q_{l}$, $p_{j},q_{j}$, or $p_{k},q_{k}$ is non-zero. Then, it must be that, for two indices (without loss of generality, $j$ and $k$, either $p_{j}>q_{j}$ and $p_{k}>q_{k}$, or $q_{j}>p_{j}$ and $q_{k}>p_{k}$. Without loss of generality, assume we are in the first case. Furthermore, without loss of generality, assume that $\frac{p_{k}}{q_{k}}\geq\frac{p_{j}}{q_{j}}$.
Now, consider our proof above where we analyze $\Pr[a=j\text{ and }b=j]$. We claim that at least one of the following inequalities from the end of the proof must be strict:
$\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}$ $\leq 1-p_{j}$ or $\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}$ $\geq p_{j}.$
Specifically, if $\frac{p_{j}}{q_{j}}=\frac{p_{k}}{q_{k}}$, then $\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}\geq p_{j}+p_{k}>p_{k}$, so the second inequality is strict. On the other hand, if $\frac{p_{k}}{q_{k}}>\frac{p_{j}}{q_{j}}$ then the first summation excludes both $p_{j},p_{k}$, which means $\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}\leq 1-p_{j}-p_{k}<1-p_{j}$.
Another example discussed in Bavarian et al. (2020) is when $\mathcal{P}$ and $\mathcal{Q}$ are both uniform, but over different sized subsets (call them $\mathcal{A}$ and $\mathcal{B}$) of $\{1,\ldots,n\}$. In this special case, it can be checked that Gumbel sampling is exactly equivalent to Broder’s “unweighted MinHash” method (Broder, 1997). It yields collision probability $\frac{|\mathcal{A}\cap\mathcal{B}|}{|\mathcal{A}\cap\mathcal{B}|}$, whereas Weighted MinHash achieves a strictly worse probability of $\frac{\left(1+\left|1/|\mathcal{A}|-1/|\mathcal{B}|\right|\right)\cdot|\mathcal{A}\cap\mathcal{B}|}{|\mathcal{A}\cap\mathcal{B}|+|\mathcal{B}|-|\mathcal{A}|}$.